1.)what is the solution to the given system of equation?
y=4x+12
3x+2y+46

2.)The lines of sight to a forest fire as follow.
from ranger station A: y=3x-11
from ranger station B: 2x+3y=0
what are the coordinates of the fire, if the fire is where the lines of sight cross?

Answer by stanbon(75887)

1.)what is the solution to the given system of equation?
y=4x+12
3x+2y+46
Comment: Where is the equal sign?
==========================================
2.)The lines of sight to a forest fire as follow.
from ranger station A: y=3x-11
from ranger station B: 2x+3y=0
what are the coordinates of the fire, if the fire is where the lines of sight cross?
---
y = 3x-11
y = (-2/3)x
----------------
3x-11 = (-2/3)x
9x-33 = -2x
11x = 33
x = 3
---
If y = (-2/3)x, y = -2
----------------------------
The fire is at (3,-2)
===========================
Cheers,
Stan H.